# A Comprehensive Evaluation and Practice of System Penetration Testing

### v1

@@ 2.2 Primary Penetration Testing Methodologies
The core purpose of this type of penetration testing is to reduce the burden on security testers while reducing labor and time costs [9]. However, it also suffers from drawbacks, such as reduced accuracy. Such attacks are typically achieved through technologies such as machine learning and deep learning. For example, Zhou et al. proposed NIG-AP, an automated penetration testing algorithm that leverages Markov decision processes and network information gain. This algorithm primarily employs reinforcement learning models and network information gain to guide the discovery of attack paths. During testing, the attacker maximizes the target network’s information entropy through a series of actions [24]. This network information entropy is comprised of two components: the host information entropy and the network information entropy. The host information entropy is further subdivided into four constituent elements: the operating system, port services, applications, and protection mechanisms. Let $P_{OS}$ denote the operating system information vector, and MM denote the other three types of information vectors. The calculation method for the target host’s exposed status information is shown in the following formula.
@@ 4.1.1 Integrated Platforms and Scanning Tools
Network Mapper (Nmap) is an open-source network discovery and security auditing tool first released by Gordon Lyon in 1997. Its stability, scalability, flexibility, and compatibility have made it highly favored by users, enabling it to stand out among numerous scanning tools and become a mainstream solution. In penetration testing, it is typically employed during the information gathering phase. It can probe not only individual IP addresses but also large IP address ranges to scan multiple hosts simultaneously. This tool enables users to accurately identify online hosts, open ports, and associated services within a network target. This information can then be leveraged to infer the types of security devices deployed in the network infrastructure, such as firewall models and filtering rules. Furthermore, it can be used in conjunction with its graphical interface, Zenmap, to conduct effective penetration testing on target systems based on the collected information.
@@
Nmap is a commonly used tool during the information gathering phase. It supports scanning across multiple protocols, including TCP, UDP, and ICMP, offers over ten scanning techniques, such as SYN, ACK, and Null, to adapt to diverse network environments, and features a continuously updated NSE script library with advanced capabilities such as service identification and vulnerability detection. Although Nmap offers exceptional precision and depth, it has some limitations. Full-port scans on large networks can be slow, which makes Nmap less efficient than Masscan. Advanced features, such as NSE scripting, require a significant learning investment. Certain scans, such as SYN, may lack stealth and could trigger enterprise IDS detection.
@@ 4.1.2 Vulnerability Management and Exploitation Tools
Nessus is a benchmark vulnerability management tool developed by Tenable Network Security, which specializes in comprehensively identifying security risks across network assets. It integrates over 20,000 vulnerabilities, covering CVE flaws, configuration errors, weak passwords, and missing patches, supporting multi-dimensional vulnerability detection. It enables real-time updates of vulnerability detection scripts with daily synchronization of the latest threat intelligence. The built-in templates for PCI, DSS, HIPAA, and other standards support compliance audits. Furthermore, customizable scanning policies allow adjustment of sensitivity before vulnerability scans. During scans, an advanced analytics engine precisely assesses risks, generating detailed risk assessment reports that categorize vulnerabilities into five severity levels: Critical, High, Medium, Low, and Info. It also integrates PDF, HTML, and SIEM systems, supporting export to multiple report formats, including PDF. Its interface is clean and intuitive, allowing quick mastery. Simply click on the scan results to view the vulnerability details, remediation recommendations, and CVSS scores. Consequently, Nessus is widely deployed for enterprise security baseline establishment and continuous threat monitoring. The drawback is its relatively high cost, with expensive commercial licensing fees making it more suitable for large enterprises. The free version limits IP scans, cannot generate customized reports, lacks enterprise-grade features, and has other core functionalities. Scanning demands significant CPU and memory resources, and some vulnerabilities rely on banner matching, potentially leading to false positives.
@@
OpenVAS is an open-source vulnerability scanning and management framework that evolved from the early Nessus source code. Positioned as a branch alternative to Nessus, it focuses on delivering enterprise-grade vulnerability detection capabilities and is primarily maintained by Greenbone Networks. OpenVAS offers capabilities including misconfiguration detection, CVE vulnerability scanning, patch management, and web application vulnerability detection. However, it heavily relies on the Network Vulnerability Tests (NVT) plugin library. It supports API integration and report export in PDF, HTML, and XML formats. While the enterprise edition requires a license fee, the community edition is fully open-source with no restrictions on scan scale, IP count, or asset quantity. It lags in vulnerability database updates compared to Nessus, which offers real-time updates and faster zero-day vulnerability coverage. In addition, OpenVAS features a complex user interface with a steep configuration learning curve, requiring significant time investment. Enterprise-level capabilities such as distributed scanning necessitate reliance on the paid Creenbone Enterprise Edition. Therefore, OpenVAS suits security teams or enterprises with limited budgets requiring large-scale asset scanning. Users must possess strong technical capabilities and be willing to invest time in configuring and maintaining open-source tools. However, its compliance requirements are relatively low, primarily focusing on basic vulnerability management.
@@
Metasploit is an open-source penetration testing framework, initially released by H.D. Moore in 2003 and currently maintained by Rapid7. Its primary purpose is to standardize the exploitation process through modular design, covering the entire attack chain from vulnerability discovery and exploitation to privilege escalation and post-exploitation. To date, Metasploit contains over 5,000 exploit modules, with vulnerability information continuously updated. In addition, the tool encompasses various functionalities from reconnaissance to reporting phases and supports multi-platform installation on Linux, Windows, and macOS, making it highly popular among cybersecurity professionals.
@@ 4.1.3 Web Application Testing Tools
SQLMap is an open-source penetration testing tool that specializes in automated detection and SQL injection. As the most representative professional tool in SQL injection testing, it focuses on a highly intelligent parameter parsing and exploit engine, widely used in database security testing. It can support mainstream injection techniques such as Boolean blind, time-based blind, error-based, and union queries. It is also compatible with major databases, such as MySQL, Oracle, and PostgreSQL. It autonomously identifies the injected parameters and pairs them with optimal attack payloads. It can directly export database table structures, field contents, and even file system or operating system-level data. In addition, it supports post-exploitation operations, such as file reading and writing, operating system command execution, and hash cracking. It also integrates with tools such as Burp Suite for log importation and enables traffic manipulation via proxies. However, it is only capable of detecting SQL injection, so it requires integration with other tools for comprehensive penetration testing. It is not suitable for covert testing and may trigger WAF alerts during automated attacks. Using os-shell to execute operating system commands may inadvertently damage the target system.
@@ 4.1.4 Automated Penetration Testing Tools
There are currently various automated penetration testing tools available, such as AutoSploit. AutoSploit is a Python-based tool that integrates search engines, such as Shodan and Quake, to identify potential attack targets. Additionally, it incorporates over 300 Metasploit attack modules and allows for the addition of new modules via configuration files. Once targets are identified, it automatically invokes these modules to execute exploit attacks. Although automated penetration testing tools alleviate the burden on security personnel, they have the following drawbacks. It updates attack payloads slowly and struggles to maintain consistent updates. Ensuring accuracy during testing is difficult, and it lacks flexibility.
@@ 4.2 Scene-Based Tool Effectiveness Evaluation Model
In actual penetration testing, the tools used vary depending on the testing objectives and specific circumstances. Generally, penetration testing tools incorporate several or all of the following functionalities: Host Scanning, Password Cracking, Web Scanning, Social Engineering, Vulnerability Discovery, Exploit, Session Control, Report Generation, and Visualization Interface. For convenience in subsequent discussions, these functions are represented in the order listed above using the initial letters of their names. Repeated letters use the initial letter of the next word, i.e., “H, P, W, S, V, E, C, R, I”. A value of 1 indicates that the tool possesses the function, while a value of 0 indicates that it does not. This paper assigns weights to these functions and uses a weighted sum formula to calculate the sum for different tools. This serves as a criterion for selecting efficient penetration testing tools. This paper primarily proposes weighting allocation schemes tailored for three distinct scenarios. Scheme 1 balances vulnerability discovery, exploitation, and baseline scanning. It is suitable for routine penetration testing, as shown in Table 7. Scheme 2 is designed for corporate intranet security assessments, as illustrated in Table 8. Scheme 3 emphasizes practical application and stealth and applies to red team attack exercises, as shown in Table 9.
@@
The above permission allocation is based on industry standards, common security threats, and their impact. Due to varying actual conditions, the weight distribution for each function may differ. Therefore, weighting must be allocated based on actual circumstances while adhering to industry standards. This ensures that penetration testing focuses on the most critical security domains and also covers other important supporting functions, thereby delivering a comprehensive security assessment. The features of commonly used cybersecurity tools are compared in Table 10. The weighted sum calculation formula is as follows.
@@ 4.3.1 Experimental Environment and Data Collection
This experiment utilizes VMware Workstation to deploy test hosts running different operating systems on a single physical machine through virtualization technology. The simulation primarily uses Kali Linux for the attacker’s host, Windows 7 for internal Windows terminal hosts within the corporate network, Windows Server 2012 R2 for internal backend server hosts, and Metasploitable 2 for internal Linux terminal hosts. Then, attacks are reproduced using the penetration testing process described in Section 3. The experimental topology is illustrated in Figure 6. The details of the virtual machines are listed in Table 11.
@@
The first step in the testing process is to gather information about the target system using active information gathering methods. This involves a direct interaction with the target system to obtain more information. Host scanning is the first step in information gathering.
@@
After obtaining the IPs of the target hosts, further information about the target hosts is required to ensure effective penetration testing. After determining the operating system information for each host in this phase, the appropriate attack modules can be selected for them. This experiment uses the Nmap tool for operating system identification. The command nmap -o [target IP] retrieves the target host’s operating system information. If the information is unclear, add the parameters -sV or -A to obtain more detailed data. Detailed information is shown in Table 12.
@@
Port scanning is a critical step in information gathering, enabling testers to understand the details of each network port fully. This allows them to infer potential attack methods based on the service types of different ports, laying the groundwork for subsequent attacks. The primary tool is Nmap, and key port scanning information is shown in Table 13.
@@
192.168.233.135 have respectively enabled the FTP remote transfer protocol, SSH remote connection, and Telnet remote connection. Therefore, attacks can be carried out through methods such as password sniffing, file transfer, or brute force attacks. Samba services are vulnerable to brute-force attacks, remote code execution, and unauthorized access. MySQL databases can be exploited through injection attacks, privilege escalation, and brute-force attacks. Apache and Tomcat can be targeted via web application vulnerabilities.
@@
After gathering basic information about target hosts through host discovery, OS identification, and port scanning, the next step is to analyze vulnerabilities present in each host system. Relevant tools are then used to scan the test targets against vulnerability databases, detecting system vulnerabilities and weaknesses. As mentioned earlier, Nessus is employed for vulnerability scanning in this experiment. Nessus is installed on Kali, requiring root privileges to start it via the command service nessusd start. Afterward, Nessus can be accessed through Kali’s built-in Firefox browser.
@@ 4.3.2 Exploitation and Attack Reproduction
Based on the results of the vulnerability scan, the host 192.168.233.131 was found to have multiple vulnerabilities of varying severity levels. This experiment selects the CVE-2019-0708 vulnerability and the MS17-010 vulnerability for attack reproduction.
@@
The MS17-010 vulnerability, also known as EternalBlue, is a set of remote code execution flaws in the SMB protocol. Attackers exploit this by sending maliciously crafted packets to port 445 to execute malicious code and gain control of the system.
@@
During the information gathering phase, the host 192.168.233.134 was found to have multiple vulnerabilities, including “VNC Server ‘password’ Password”, “Apache Tomcat AJP Connector Request Injection”, and “Samba Badlock Vulnerability”. This experiment primarily exploits these three vulnerabilities for attacks.
@@
192.168.233.134 to configure the target host. Next, run show payload to view the available payloads. Finally, set the payload with set payload cmd/unix/reverse. Execute with run to successfully launch the attack.
@@ 4.4.1 File Upload Vulnerabilities
File upload and download are indispensable functions in web applications, designed to fulfill legitimate user needs such as uploading images or documents. However, if servers are poorly designed and fail to rigorously filter uploaded files, this critical feature can be exploited by malicious actors to upload malicious scripts, ultimately gaining server privileges. Attackers typically upload web backdoor files, also known as Webshells, commonly written in PHP, ASP, or JSP. They bypass security checks primarily through methods such as GET form manipulation. Once successfully uploaded, these backdoors enable remote command execution via web access. This attack is persistent—once established, attackers can maintain long-term control over the web server.
In DVWA, the “low” security level disables file type and size checks, allowing arbitrary file uploads. However, the “medium” mode imposes restrictions on file types and sizes, preventing the successful upload of PHP files. Based on the error message, only JPEG or PNG image files are permitted. To successfully upload a file, open Burp Suite, switch to the Proxy tab, and use the built-in browser to access DVWA via “Open Browser”. Then click “Intercept” to capture HTTP packets. Change the “Content-Type” to “image/jpeg”, and send the data packet to bypass detection and successfully upload the file.

### v2

@@ 2.2 Primary Penetration Testing Methodologies
The core purpose of this type of penetration testing is to reduce the burden on security testers while reducing labor and time costs [18]. However, it also suffers from drawbacks, such as reduced accuracy. Such attacks are typically achieved through technologies such as machine learning and deep learning. For example, Zhou et al. proposed NIG-AP, an automated penetration testing algorithm that leverages Markov decision processes and network information gain. This algorithm primarily employs reinforcement learning models and network information gain to guide the discovery of attack paths. During testing, the attacker maximizes the target network’s information entropy through a series of actions [41]. This network information entropy is comprised of two components: the host information entropy and the network information entropy. The host information entropy is further subdivided into four constituent elements: the operating system, port services, applications, and protection mechanisms. Let $P_{OS}$ denote the operating system information vector, and $M$ denote the other three types of information vectors. The calculation method for the target host’s exposed status information is shown in the following formula.
@@ 4.1.1 Integrated Platforms and Scanning Tools
Network Mapper (Nmap) is an open-source network discovery and security auditing tool first released by Gordon Lyon in 1997. Its stability, scalability, flexibility, and compatibility have made it highly favored by users, enabling it to stand out among numerous scanning tools and become a mainstream solution [1]. In penetration testing, it is typically employed during the information gathering phase. It can probe not only individual IP addresses but also large IP address ranges to scan multiple hosts simultaneously. This tool enables users to accurately identify online hosts, open ports, and associated services within a network target. This information can then be leveraged to infer the types of security devices deployed in the network infrastructure, such as firewall models and filtering rules. Furthermore, it can be used in conjunction with its graphical interface, Zenmap, to conduct effective penetration testing on target systems based on the collected information.
@@
Nmap is a commonly used tool during the information gathering phase [48]. It supports scanning across multiple protocols, including TCP, UDP, and ICMP, offers over ten scanning techniques, such as SYN, ACK, and Null, to adapt to diverse network environments, and features a continuously updated NSE script library with advanced capabilities such as service identification and vulnerability detection. Although Nmap offers exceptional precision and depth, it has some limitations. Full-port scans on large networks can be slow, which makes Nmap less efficient than Masscan. Advanced features, such as NSE scripting, require a significant learning investment. Certain scans, such as SYN, may lack stealth and could trigger enterprise IDS detection.
@@ 4.1.2 Vulnerability Management and Exploitation Tools
Nessus is a benchmark vulnerability management tool developed by Tenable Network Security, which specializes in comprehensively identifying security risks across network assets. It integrates over 20,000 vulnerabilities, covering CVE flaws, configuration errors, weak passwords, and missing patches, supporting multi-dimensional vulnerability detection [32]. It enables real-time updates of vulnerability detection scripts with daily synchronization of the latest threat intelligence. The built-in templates for PCI, DSS, HIPAA, and other standards support compliance audits. Furthermore, customizable scanning policies allow adjustment of sensitivity before vulnerability scans. During scans, an advanced analytics engine precisely assesses risks, generating detailed risk assessment reports that categorize vulnerabilities into five severity levels: Critical, High, Medium, Low, and Info. It also integrates PDF, HTML, and SIEM systems, supporting export to multiple report formats, including PDF. Its interface is clean and intuitive, allowing quick mastery. Simply click on the scan results to view the vulnerability details, remediation recommendations, and CVSS scores. Consequently, Nessus is widely deployed for enterprise security baseline establishment and continuous threat monitoring. The drawback is its relatively high cost, with expensive commercial licensing fees making it more suitable for large enterprises. The free version limits IP scans, cannot generate customized reports, lacks enterprise-grade features, and has other core functionalities. Scanning demands significant CPU and memory resources, and some vulnerabilities rely on banner matching, potentially leading to false positives.
@@
OpenVAS is an open-source vulnerability scanning and management framework that evolved from the early Nessus source code. Positioned as a branch alternative to Nessus, it focuses on delivering enterprise-grade vulnerability detection capabilities and is primarily maintained by Greenbone Networks. OpenVAS offers capabilities including misconfiguration detection, CVE vulnerability scanning, patch management, and web application vulnerability detection. However, it heavily relies on the Network Vulnerability Tests (NVT) plugin library. It supports API integration and report export in PDF, HTML, and XML formats. While the enterprise edition requires a license fee, the community edition is fully open-source with no restrictions on scan scale, IP count, or asset quantity [43]. It lags in vulnerability database updates compared to Nessus, which offers real-time updates and faster zero-day vulnerability coverage. In addition, OpenVAS features a complex user interface with a steep configuration learning curve, requiring significant time investment. Enterprise-level capabilities such as distributed scanning necessitate reliance on the paid Creenbone Enterprise Edition. Therefore, OpenVAS suits security teams or enterprises with limited budgets requiring large-scale asset scanning. Users must possess strong technical capabilities and be willing to invest time in configuring and maintaining open-source tools. However, its compliance requirements are relatively low, primarily focusing on basic vulnerability management.
@@
Metasploit is an open-source penetration testing framework, initially released by H.D. Moore in 2003 and currently maintained by Rapid7 [37]. Its primary purpose is to standardize the exploitation process through modular design, covering the entire attack chain from vulnerability discovery and exploitation to privilege escalation and post-exploitation. To date, Metasploit contains over 5,000 exploit modules, with vulnerability information continuously updated. In addition, the tool encompasses various functionalities from reconnaissance to reporting phases and supports multi-platform installation on Linux, Windows, and macOS, making it highly popular among cybersecurity professionals.
@@ 4.1.3 Web Application Testing Tools
SQLMap is an open-source penetration testing tool that specializes in automated detection and SQL injection. As the most representative professional tool in SQL injection testing, it focuses on a highly intelligent parameter parsing and exploit engine, widely used in database security testing. It can support mainstream injection techniques such as Boolean blind, time-based blind, error-based, and union queries [51]. It is also compatible with major databases, such as MySQL, Oracle, and PostgreSQL. It autonomously identifies the injected parameters and pairs them with optimal attack payloads. It can directly export database table structures, field contents, and even file system or operating system-level data. In addition, it supports post-exploitation operations, such as file reading and writing, operating system command execution, and hash cracking. It also integrates with tools such as Burp Suite for log importation and enables traffic manipulation via proxies. However, it is only capable of detecting SQL injection, so it requires integration with other tools for comprehensive penetration testing. It is not suitable for covert testing and may trigger WAF alerts during automated attacks. Using os-shell to execute operating system commands may inadvertently damage the target system.
@@ 4.1.4 Automated Penetration Testing Tools
There are currently various automated penetration testing tools available, such as AutoSploit. AutoSploit is a Python-based tool that integrates search engines, such as Shodan and Quake, to identify potential attack targets [15]. Additionally, it incorporates over 300 Metasploit attack modules and allows for the addition of new modules via configuration files. Once targets are identified, it automatically invokes these modules to execute exploit attacks. Although automated penetration testing tools alleviate the burden on security personnel, they have the following drawbacks. It updates attack payloads slowly and struggles to maintain consistent updates. Ensuring accuracy during testing is difficult, and it lacks flexibility.
@@ 4.2 Scene-Based Tool Effectiveness Evaluation Model
In actual penetration testing, the tools used vary depending on the testing objectives and specific circumstances. Generally, penetration testing tools incorporate several or all of the following functionalities: Host Scanning, Password Cracking, Web Scanning, Social Engineering, Vulnerability Discovery, Exploit, Session Control, Report Generation, and Visualization Interface [13]. For convenience in subsequent discussions, these functions are represented in the order listed above using the initial letters of their names. Repeated letters use the initial letter of the next word, i.e., “H, P, W, S, V, E, C, R, I”. A value of 1 indicates that the tool possesses the function, while a value of 0 indicates that it does not. This paper assigns weights to these functions and uses a weighted sum formula to calculate the sum for different tools. This serves as a criterion for selecting efficient penetration testing tools. This paper primarily proposes weighting allocation schemes tailored for three distinct scenarios. Scheme 1 balances vulnerability discovery, exploitation, and baseline scanning. It is suitable for routine penetration testing, as shown in Table 7. Scheme 2 is designed for corporate intranet security assessments, as illustrated in Table 8. Scheme 3 emphasizes practical application and stealth and applies to red team attack exercises, as shown in Table 9.
@@
The above permission allocation is based on industry standards, common security threats, and their impact. Due to varying actual conditions, the weight distribution for each function may differ [38]. Therefore, weighting must be allocated based on actual circumstances while adhering to industry standards. This ensures that penetration testing focuses on the most critical security domains and also covers other important supporting functions, thereby delivering a comprehensive security assessment. The features of commonly used cybersecurity tools are compared in Table 10. The weighted sum calculation formula is as follows.
@@ 4.3.1 Experimental Environment and Data Collection
This experiment utilizes VMware Workstation to deploy test hosts running different operating systems on a single physical machine through virtualization technology [28]. The simulation primarily uses Kali Linux for the attacker’s host, Windows 7 for internal Windows terminal hosts within the corporate network, Windows Server 2012 R2 for internal backend server hosts, and Metasploitable 2 for internal Linux terminal hosts. Then, attacks are reproduced using the penetration testing process described in Section 3. The experimental topology is illustrated in Figure 6. The details of the virtual machines are listed in Table 11.
@@
The first step in the testing process is to gather information about the target system using active information gathering methods [46]. This involves a direct interaction with the target system to obtain more information. Host scanning is the first step in information gathering.
@@
After obtaining the IPs of the target hosts, further information about the target hosts is required to ensure effective penetration testing. After determining the operating system information for each host in this phase, the appropriate attack modules can be selected for them [45]. This experiment uses the Nmap tool for operating system identification. The command nmap -o [target IP] retrieves the target host’s operating system information. If the information is unclear, add the parameters -sV or -A to obtain more detailed data. Detailed information is shown in Table 12.
@@
Port scanning is a critical step in information gathering, enabling testers to understand the details of each network port fully [42]. This allows them to infer potential attack methods based on the service types of different ports, laying the groundwork for subsequent attacks. The primary tool is Nmap, and key port scanning information is shown in Table 13.
@@
192.168.233.135 have respectively enabled the FTP remote transfer protocol, SSH remote connection, and Telnet remote connection. Therefore, attacks can be carried out through methods such as password sniffing, file transfer, or brute force attacks [3]. Samba services are vulnerable to brute-force attacks, remote code execution, and unauthorized access. MySQL databases can be exploited through injection attacks, privilege escalation, and brute-force attacks. Apache and Tomcat can be targeted via web application vulnerabilities.
@@
After gathering basic information about target hosts through host discovery, OS identification, and port scanning, the next step is to analyze vulnerabilities present in each host system. Relevant tools are then used to scan the test targets against vulnerability databases, detecting system vulnerabilities and weaknesses. As mentioned earlier, Nessus is employed for vulnerability scanning in this experiment. Nessus is installed on Kali, requiring root privileges to start it via the command service nessusd start. Afterward, Nessus can be accessed through Kali’s built-in Firefox browser [9].
@@ 4.3.2 Exploitation and Attack Reproduction
Based on the results of the vulnerability scan, the host 192.168.233.131 was found to have multiple vulnerabilities of varying severity levels [2]. This experiment selects the CVE-2019-0708 vulnerability and the MS17-010 vulnerability for attack reproduction.
@@
The MS17-010 vulnerability, also known as EternalBlue, is a set of remote code execution flaws in the SMB protocol. Attackers exploit this by sending maliciously crafted packets to port 445 to execute malicious code and gain control of the system [29].
@@
During the information gathering phase, the host 192.168.233.134 was found to have multiple vulnerabilities, including “VNC Server ‘password’ Password”, “Apache Tomcat AJP Connector Request Injection”, and “Samba Badlock Vulnerability” [52]. This experiment primarily exploits these three vulnerabilities for attacks.
@@
192.168.233.134 to configure the target host. Next, run show payload to view the available payloads. Finally, set the payload with set payload cmd/unix/reverse. Execute with run to successfully launch the attack [44].
@@ 4.4.1 File Upload Vulnerabilities
File upload and download are indispensable functions in web applications, designed to fulfill legitimate user needs such as uploading images or documents [53]. However, if servers are poorly designed and fail to rigorously filter uploaded files, this critical feature can be exploited by malicious actors to upload malicious scripts, ultimately gaining server privileges. Attackers typically upload web backdoor files, also known as Webshells, commonly written in PHP, ASP, or JSP. They bypass security checks primarily through methods such as GET form manipulation. Once successfully uploaded, these backdoors enable remote command execution via web access. This attack is persistent—once established, attackers can maintain long-term control over the web server.
In DVWA, the “low” security level disables file type and size checks, allowing arbitrary file uploads [19]. However, the “medium” mode imposes restrictions on file types and sizes, preventing the successful upload of PHP files. Based on the error message, only JPEG or PNG image files are permitted. To successfully upload a file, open Burp Suite, switch to the Proxy tab, and use the built-in browser to access DVWA via “Open Browser”. Then click “Intercept” to capture HTTP packets. Change the “Content-Type” to “image/jpeg”, and send the data packet to bypass detection and successfully upload the file.